The kinetic energy of a cougar that was moving at a speed of 35km/h was determined to be 4000J. What was the mass of the cougar?
1
Expert's answer
2013-01-23T10:55:28-0500
The kinetic energy of body of mass m moving at speed v isgiven by the formula
E = m v^2 / 2,
whencem = 2E / v^2.

In our case
m = 2 * 4000 / 35^2 = 8000 / 1225 = 6.53 kg